an ancient manuscript reveals that a landowner in the land of Middle Earth held 3 arces of plowed land plus a livestock area of 23 perches by 3 perches. One perch is 16.5 ft. One arce is 40 perches by 4 perches.

1-What is the area of the landowner in arces? Aarces=

2-What is the area of the landowner in roods? (1 rood is 40 perches by 1 perch.) Aroods
Answer #1

A. 1 arce = 160 perch^2
Land area = 3+(23*3)/160 arces=3.431 acres
B. 1 rood = 40 perch^2 = 0.25 arce
Land area = (3+(23*3)/160)/0.25 roods= 13.724 roods